The length of time patients must wait to see a doctor at an emergency room in a large hospital is uniformly distributed between 40 minutes and 3 hours.
a) What is the probability that a patient would have to wait between 50 minutes and two hours?
b) What is the probability that a patient would have to wait exactly one hour?
c) What is the probability that a patient would have to wait between 60 and 90 minutes?
d) What is the probability that a patient would have to wait between 120 minutes and 3 hours?
e) What is the probability that a patient would have to wait between 60 and 120 minutes?

Answer by stanbon(29603)

The length of time patients must wait to see a doctor at an emergency room in a large hospital is uniformly distributed between 40 minutes and 3 hours.
a) What is the probability that a patient would have to wait between 50 minutes and two hours?
(120-50)140=70/180=7/18
b) What is the probability that a patient would have to wait exactly one hour?
60/180=1/3
c) What is the probability that a patient would have to wait between 60 and 90 minutes?
(90-60)/180=30/180=1/3
d) What is the probability that a patient would have to wait between 120 minutes and 3 hours?
(180-120)/180=1/3
e) What is the probability that a patient would have to wait between 60 and 120 minutes?
(120-60)/180=1/3
-----------
Cheers,
Stan H.